Enterprise Group issued $100,000 of 4-year, 6% bonds outstanding on December 31, 2015 for $92,000. Enterprise uses straight-line amortization. On April 1, 2016, $50,000 of the bonds were retired at 97. What is the book value of the bonds sold on April 1?
Answer #1

Solution:

Face value of bonds= $100,000,

Issue price of bonds= $92,000,

Discount on bonds = $100,000 - $92,000,

= $8,000,

Percentage of bonds retired = 50,000/100,000 = 0.5,

So discount on 8,000*0.5 = $4,000

Amortization of bonds $4,000/ 48 = 83.33 per month,

So on 1st April'2016 discount Amortized = 83 * 3,

= $249,

Book value of bonds on April 1,2016

Bonds Payable =$50,000

less:Discount on bonds = $ 4,000 - $249,

=$50,000 - $ 3,751,

=$46,249

Book value of bonds on April 1,2016 $46,249.

Please give a Positive Rating...